The length of a rectangle is 1 inch more than four times its width. The perimeter is 52 inches. Find the length and width of the rectangle.
Answer by Nada(11)

l > is the length & w > is the width
Because a perimater of a rectangle is > 2( Width + Length )
& w = ( w ) & l = ( 4w+1 )
then the equation will be 52 = 2 ( (w) + (4w+1) )
52 = 2w+8w+2
52 = 10w+2
10w = 52-2 >> w = 50/10 = 2 in
Then to get the length :
52 = 2(5+L)
52 = 10+2L
2L = 52-10 >> L = 42/2 = 21
To check your answer :
2(5+21) = 52 ( answer is correct)